OpenStudy (anonymous):

Estimate the number of breaths taken during 70 years, the average life span of a person

OpenStudy (anonymous):

how many breaths you take per minute x \[ \frac{ 1 hr }{ 60 minutes } x \frac{ 1day }{ 24 hours } x \frac{ 1 year }{ 365 days } x \frac{ Avg lifespan }{ 70 years }\]